OpenStudy (anonymous):

HELP!! WILL GIVE MEDAL Solve -x/4 ≤ 7

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

I got greater than or equal to -28

sammixboo (sammixboo):

Okie dokie, so all you need to do is simplify the problem to look like this \(\rm \color{#20bd23}{\dfrac{-1}{4}x=~7}\) Now divide \(\rm \color{#20bd23}{\dfrac{-1}{4}}\) or -0.25 on both sides of the equation \(\rm \color{#20bd23}{\dfrac{-0.25x}{-0.25}=\dfrac{7}{-0.25}}\) \(\rm \color{#20bd23}{x=\dfrac{7}{-0.25}}\) What is 7/-0/25?

sammixboo (sammixboo):

And that is correct :)

OpenStudy (anonymous):

Thanks can u check another?

OpenStudy (anonymous):

The equation below represents the fact that your driving rate (the distance divided by the time in hours) must be less than or equal to 65 mph by law. If you drove for 2.5 hours, what is the maximum distance you can drive and still be within the law? Express your answer as an inequality with x as the variable. \[\frac{ distance }{ time}\le 65\]

OpenStudy (anonymous):

I got x greater than or equal to 162.5

sammixboo (sammixboo):

Well, if you drove for 2.5 hours, that is our time \(\rm \color{#20bd23}{\dfrac{x}{2.5}\le~65}\) All you need to do is multiply 65 x 2.5, which is 162.5

OpenStudy (anonymous):

Cool. Ty so much :)